The Mount Emerald Wind Farm will be Queenslands largest wind farm once construction is completed. Nearby cattle farmers complain that the wind farm will adversely affect their cattles wellbeing. Energex is the state power producer who will operate the wind farm. No Wind Farm Wind Farm 纉) 《.ialt, tポ看ii.T.($) Gains to Energex (S Total (S 10 210 Consider Energex had o payout all the extra gans he received rom construction o the wind armsto cattle armers if a nego a on costs a er were 201 and armers were affected, how much would each farmer receive in compensation if paid out equally? Answer to the nearest whole number in dollars (with no decimal places, S sign, spaces or commas)
Answer #1

When there is no wind farm, the gains to enengex is $10. After the wind farm is built , the gains to energex becomes $600. So the total extra gains Energex received is 600-10 = $590

Negotiation costs with the lawyers = 231

The total gains after considering the negotiation costs = 590-231 = 359

And now this gains needs to be distributed equally among 6 farmers.

So the share of each farmer will be = 359/6 = 59.833

Rounding off to nearest whole number, we get the amount to be 60.

So each farmer will get 60 as compensation.